# Adaptive Online Bayesian Estimation of Frequency Distributions with Local Differential Privacy

### v1

@@ Abstract
We propose a novel Bayesian approach for the adaptive and online estimation of the frequency distribution of a finite number of categories under the local differential privacy (LDP) framework. The proposed algorithm performs Bayesian parameter estimation via posterior sampling and adapts the randomization mechanism for LDP based on the obtained posterior samples. We propose a randomized mechanism for LDP which uses a subset of categories as an input and whose performance depends on the selected subset and the true frequency distribution. By using the posterior sample as an estimate of the frequency distribution, the algorithm performs a computationally tractable subset selection step to maximize the utility of the privatized response of the next user. We propose several utility functions related to well-known information metrics, such as (but not limited to) Fisher information matrix, total variation distance, and information entropy. We compare each of these utility metrics in terms of their computational complexity. We employ stochastic gradient Langevin dynamics for posterior sampling, a computationally efficient approximate Markov chain Monte Carlo method. We provide a theoretical analysis showing that (i) the posterior distribution targeted by the algorithm converges to the true parameter even for approximate posterior sampling, and (ii) the algorithm selects the optimal subset with high probability if posterior sampling is performed exactly. We also provide numerical results that empirically demonstrate the estimation accuracy of our algorithm where we compare it with nonadaptive and semi-adaptive approaches under experimental settings with various combinations of privacy parameters and population distribution parameters.
@@ 1 Introduction
Frequency distributions are encountered in many applications that involve personal and private data. Suppose a type of sensitive information is represented as a categorical random variable $X$ which has a population distribution denoted by $\text{Cat}(\theta)$, where $\theta$ is a $K$-dimensional probability vector.
In this paper, we propose an adaptive and online algorithm to estimate $\theta$ in a Local Differential Privacy (LDP) framework, when $X$ is unobserved and instead, we have access to a randomized response $Y$ derived from $X$. In the LDP framework, a central aggregator receives privatized/perturbed data from each user and uses this noisy data for inferential tasks. In that sense, LDP differs from standard DP framework (Dwork, , 2006) where the aggregator privatizes operations on the sensitive dataset after it collects the sensitive data without noise. Hence LDP can be said to provide a stricter form of privacy and is used in cases where the aggregator may not be trustable (Kasiviswanathan et al., , 2011). This feature of LDP makes it especially useful for users of social media platforms and cloud computing services where the central aggregator is more prone to leakage of personal information with or without the knowledge of the company that aggregates the data. Below we give a more formal definition of $\epsilon$-LDP as a property that concerns a randomized mechanism.
@@
The definition of LDP is almost the same as that of global DP. The main difference is that, in the global DP, neighboring inputs $x,x^{\prime}$ are two datasets that differ in only one record; whereas in LDP, $x,x^{\prime}$ pairs are two different data points from the same domain. In Definition 1, $\epsilon\geq 0$ is the privacy parameter. A smaller $\epsilon$ value provides stronger privacy.
One main challenge in most differential privacy settings is to decide on the randomized mechanism. In the case of LDP, this is how an individual data point $X$ should be randomized. For a given randomized algorithm, too little randomization may not guarantee the privacy of individual users; whereas too severe randomization deteriorates the utility of the output of the randomized algorithm. Balancing these conflicting objectives (privacy vs utility) is the main goal of the research on estimation under privacy constraints.
While sequential observation of the individual points in a dynamic/online setting may make the estimation task under the LDP constraint harder, it may offer an opportunity to adapt the randomized mechanism in time to improve the estimation quality. Motivated by that, in this paper, we address the problem of online Bayesian estimation of a frequency distribution ($\theta$) under $\epsilon$-LDP, while at the same time choosing the randomization mechanism adaptively so that the utility is improved continually in time.
@@
This paper presents AdOBEst-LDP: a novel methodological framework for Adaptive Online Bayesian Estimation with LDP. A flowchart diagram of AdOBEst-LDP is given in Figure 1 to expose the reader to the main idea of the framework. AdOBEst-LDP seeks to adapt its randomized response mechanism to past data to generate the ‘most informative’ responses about $\theta$ while maintaining $\epsilon$-LDP for all users. AdOBEst-LDP uses a randomized mechanism, called Randomly Restricted Randomized Response (RRRR), to generate randomized responses under $\epsilon$-LDP. The RRRR mechanism is a generalization of the standard randomized response mechanism in that it restricts the response to a subset of the sample space of $X$. The subset used in RRRR is decided such that the sensitive information $X$ from the next individual is likely contained in that subset. To ensure this, the subset selection step uses two inputs: (i) a sample for $\theta$ drawn from the posterior distribution of $\theta$ conditional on the past data, (ii) a utility function that scores the informativeness of the randomized response obtained from RRRR when it is run with a given subset. To that end, we propose several utility functions to score the informativeness of the randomized response. The utility functions are based on well-known tools from probability and statistics. The first utility function is based on the Fisher information of the privatized response (Alparslan and Yıldırım, , 2022; Lopuhaä-Zwakenberg et al., , 2022; Steinberger, , 2024; Yıldırım, , 2024). We propose a second utility function based on the entropy of the randomized response. We also propose two utility functions based on total variation distance. A fifth utility function is proposed based on the expected squared error of the Bayesian estimator of the sensitive data given the randomized response. Finally, the last utility function considered is the probability of honest response, i.e., $Y=X$. We provide some insight into those utility functions both theoretically and numerically. Moreover, we also provide a computational complexity analysis for the proposed utility functions.
Figure 1: Flowchart diagram of AdOBest-LDP
@@
We equip AdOBEst-LDP with an online posterior sampling for parameter estimation. Standard MCMC methods, such as Gibbs sampling, have a computation complexity quadratic in the number of individuals whose data have been collected. As a remedy to this, similar to Mazumdar et al., (2020) we propose a stochastic gradient Langevin dynamics (SGLD)-based algorithm to obtain approximate posterior samples (Welling and Teh, , 2011). By working on subsets of data, the SGLD scales in time.
@@
The numerical experiments show that AdOBEst-LPD outperforms its non-adaptive counterpart when run with SGLD for posterior sampling. The results also suggest that the utility functions considered in this paper are robust and perform well. The MATLAB code at https://github.com/soneraydin/AdOBEst_LDP can be used to reproduce the results obtained in this paper.
@@
Finally, we provide a theoretical analysis of AdOBEst-LDP. We show two key results:
@@
    If perfect sampling is used for the posterior sampling stage, the expected frequency of the algorithm choosing the best subset (according to the utility function) tends to $1$ as $n$ goes to $\infty$.
The theoretical results depend on fairly weak, realistic, and verifiable assumptions.
@@ 2 Related Literature
Estimating frequency distributions under the LDP setting has been an increasingly popular research area in recent years. Along with its basic application (estimation of discrete probabilities from locally privatized data), it is also used for a wide range of other estimation and learning purposes; to name a few, estimation of confidence intervals and confidence sets for a population mean (Waudby-Smith et al., , 2023), estimation or identification of heavy hitters (Zhu et al., , 2024) (Wang et al., 2024b, ) (Jia and Gong, , 2019), estimation of quantiles (Cormode and Bharadwaj, , 2022), frequent itemset mining (Zhao et al., , 2023), estimation of degree distribution in social networks (Wang et al., 2024a, ), distributed training of graph neural networks with categorical features and labels (Bhaila et al., , 2024). The methods that are proposed for $\epsilon$-LDP frequency estimation also form the basis of more complex inferential tasks (with some modifications on these methods), such as the release of ‘marginals’ (contingency tables) between multiple categorical features and their correlations, as in the work of Cormode et al., (2018).
@@
AdOBEst-LDP employs a modified version of the Generalized Randomized Response mechanism (GRR) to generate the randomized responses. Known also as $k$-randomized response and direct encoding, this LDP mechanism (or protocol) is the generalization of the binary randomized response mechanism for the privatization of categorical data when the number of categories is $K>2$. The basic version of GRR is based on simply flipping the category of a given response with probability $\frac{1}{e^{\epsilon}+K-1}$ or reporting its true category with probability $\frac{e^{\epsilon}}{e^{\epsilon}+K-1}$. This is a well-studied mechanism in the DP literature, and the statistical properties of its basic version (such as its estimation variance) can be found in the works by (Wang et al., , 2017) and (Wang et al., , 2020). For convenience and unity of terminology, we will call this basic approach as standard randomized response (SRR) mechanism throughout the paper. Unlike the basic version of SRR, the perturbation probability of responses in our algorithm changes adaptively, depending on the cardinality of the selected subset of categories (which we explain in detail in Section 4) for the privatization of real $X$, and the cardinality of its complementary set.
@@
The use of information metrics as utility functions in LDP protocols has been an active line of research in recent years. In the work of Kairouz et al., (2016), information metrics like $f$-divergence and mutual information are used for selecting optimal LDP protocols. In the same vein, Steinberger, (2024) uses Fisher Information as the utility metric for finding a nearly optimal LDP protocol for the frequency estimation problem, and Lopuhaä-Zwakenberg et al., (2022) uses it for comparing the utility of various LDP protocols for estimation of frequency distributions and finding the optimal one. In these works, the mentioned information metrics are used statically, i.e., to choose a protocol once and for all, for a given estimation task. The approaches in these works suffer from computational complexity for large values of $K$, because the search space for optimal protocols there grows in the order of $2^{K}$. In some other works, such as Wang et al., (2016), a randomly sampled subset of size $k\leq K$ is used to improve the efficiency of this task, where the optimal $k$ is determined by maximizing the mutual information between real data and the privatized data. However, this approach is also static, the optimal subset size $k$ is selected only once, and the optimization procedure only determines $k$ and not the subset itself. Unlike those static approaches, AdOBEst-LDP dynamically encompasses the information metric (such as the Fisher Information matrix and the other alternatives in Section 4.2) to select the optimal subset for each time step. In addition, in the subset step of AdOBEst-LDP, only $K$ candidate subsets are compared in terms of their utilities at each iteration, enabling computational traceability. This way of tackling the problem requires computing the given information metric for only $K$ times at each iteration. We will provide further details of this approach and elaborate on its computational advantages in Section 4.2.
@@
Another use of the Fisher Information in the LDP literature is for bounding the estimation error for a given LDP protocol. For example, Barnes et al., (2020) uses Fisher Information inside van Trees inequality, the Bayesian version of the Cramér-Rao bound (Gill and Levit, , 1995), for bounding the estimation error of various LDP protocols for Gaussian mean estimation and frequency estimation. Again, their work provides rules for choosing optimal protocols for a given $\epsilon$ in a static way. As a similar example, Acharya et al., (2023) derives a general information contraction bound for parameter estimation problems under LDP and shows its relation to van Trees inequality as its special case. To our knowledge, our approach is the first one that adaptively uses a utility metric to dynamically update the inner workings of an LDP protocol for estimating frequency distributions.
@@
The idea of building adaptive mechanisms for improved estimation under the LDP has been studied in the literature, although the focus and methodology of those works differ from ours. For example, Joseph et al., (2019) proposed a two-step adaptive method to estimate the unknown mean parameter of data from Gaussian distribution. In this method, the users are split into two groups, an initial mean estimate is obtained from the perturbed data of the first group, and the data from the second group is transformed adaptively according to that initial estimate. Similarly, Wei et al., (2024) proposed another two-step adaptive method for the mean estimation problem, in which, the aggregator first computes a rough distribution estimate from the noisy data of a small sample of users, which is then used for adaptively adjusting the amount of perturbation for the data of remaining users. While Joseph et al., (2019); Wei et al., (2024) consider a two-stage method, AdOBEst-LDP seeks to adapt continually by updating its LDP mechanism each time an individual’s information is collected. Similar to our work, Yıldırım, (2024) has recently proposed an adaptive LDP mechanism for online parameter estimation for continuous distributions. The LDP mechanism of Yıldırım, (2024) contains a truncation step with boundaries adapted to the estimate from the past data according to a utility function based on the Fisher information. Unfortunately, the parameter estimation step of Yıldırım, (2024) does not scale well in time. Differently from Yıldırım, (2024), AdOBEst-LDP focuses on frequency distributions, considers several other utility functions to update its LDP mechanism, employs a scalable parameter estimation step, and its performance is backed up with theoretical results.
To avoid potential confusion, we remark that other notions of ‘adaptiveness’ are also used in the DP literature. For example, Du et al., (2021) used an adaptive scheme to partition the domain of non-categorical features for a given dataset, to frugally use the privacy budget for responding to range queries. Similarly, Qardaji et al., (2013) proposed an adaptive method to partition geospatial data into grids, to answer range queries. As another example of using a different notion of adaptiveness, da Costa Filho and Machado, (2023) proposed an adaptive selection scheme to choose one of the two well-known frequency estimation protocols, namely GRR and OLH (optimal local hash), as a step of their main algorithm. Similarly, Zhang et al., (2018) proposed an adaptive scheme to choose between GRR and OUE (optimized unary encoding) protocols for a given task of marginal release. Li and Miklau, (2012) proposed an adaptive strategy to distribute the tasks of answering batch queries (which they call workload) in $(\epsilon,\delta)$-DP setting, to use a given privacy budget effectively. McKenna et al., (2020) also uses the same notion of adaptiveness for answering batch queries, but under the $\epsilon$-LDP setting.

### v2

# Bayesian Frequency Estimation Under Local Differential Privacy With an Adaptive Randomized Response Mechanism

@@ Abstract
Frequency estimation plays a critical role in many applications involving personal and private categorical data. Such data are often collected sequentially over time, making it valuable to estimate their distribution online while preserving privacy. We propose AdOBEst-LDP, a new algorithm for adaptive, online Bayesian estimation of categorical distributions under local differential privacy (LDP). The key idea behind AdOBEst-LDP is to enhance the utility of future privatized categorical data by leveraging inference from previously collected privatized data. To achieve this, AdOBEst-LDP uses a new adaptive LDP mechanism to collect privatized data. This LDP mechanism constrains its output to a subset of categories that ‘predicts’ the next user’s data. By adapting the subset selection process to the past privatized data via Bayesian estimation, the algorithm improves the utility of future privatized data. To quantify utility, we explore various well-known information metrics, including (but not limited to) the Fisher information matrix, total variation distance, and information entropy. For Bayesian estimation, we utilize posterior sampling through stochastic gradient Langevin dynamics, a computationally efficient approximate Markov chain Monte Carlo (MCMC) method.
We provide a theoretical analysis showing that (i) the posterior distribution of the category probabilities targeted with Bayesian estimation converges to the true probabilities even for approximate posterior sampling, and (ii) AdOBEst-LDP eventually selects the optimal subset for its LDP mechanism with high probability if posterior sampling is performed exactly. We also present numerical results to validate the estimation accuracy of AdOBEst-LDP. Our comparisons show its superior performance against non-adaptive and semi-adaptive competitors across different privacy levels and distributional parameters.
@@ 1 Introduction
Frequency estimation is the focus of many applications that involve personal and private categorical data. Suppose a type of sensitive information is represented as a random variable $X$ with a categorical distribution denoted by $\text{Cat}(\theta)$, where $\theta$ is a $K$-dimensional probability vector. As real-life examples, this could be the distribution of the types of a product bought by the customers of an online shopping company, responses to a poll question like “Which party will you vote for in the next elections?”, occupational affiliations of the people who visit the website of a governmental agency, and so on.
In this paper, we propose an adaptive and online algorithm to estimate $\theta$ in a Local Differential Privacy (LDP) framework where $X$ is unobserved and instead, we have access to a randomized response $Y$ derived from $X$. In the LDP framework, a central aggregator receives each user’s randomized (privatized) data to be used for inferential tasks. In that sense, LDP differs from global DP (Dwork, , 2006) where the aggregator privatizes operations on the sensitive dataset after it collects the sensitive data without noise. Hence LDP can be said to provide a stricter form of privacy and is used in cases where the aggregator may not be trustable (Kasiviswanathan et al., , 2011). Below, we give a more formal definition of $\epsilon$-LDP as a property that concerns a randomized mechanism.
@@
The definition of LDP is almost the same as that of global DP. The main difference is that, in the global DP, inputs $x,x^{\prime}$ are two datasets that differ in only one individual’s record, whereas in LDP, $x,x^{\prime}$ are two different data points from $\mathcal{X}$.
In Definition 1, $\epsilon\geq 0$ is the privacy parameter. A smaller $\epsilon$ value provides stronger privacy. One main challenge in most differential privacy settings is to decide on the randomized mechanism. In the case of LDP, this is how an individual data point $X$ should be randomized. For a given randomized algorithm, too little randomization may not guarantee the privacy of individuals, whereas too severe randomization deteriorates the utility of the output of the randomized algorithm. Balancing these conflicting objectives (privacy vs utility) is the main goal of the research on estimation under privacy constraints.
In many cases, individuals’ data points are collected sequentially. A basic example is opinion polling, where data is collected typically in time intervals of lengths in the order of hours or days. Personal data entered during registration is another example. For example, a hospital can collect patients’ categorical data as they visit the hospital for the first time.
While sequential collection of individual data may make the estimation task under the LDP constraint harder, it may also offer an opportunity to adapt the randomized mechanism in time to improve the estimation quality. Motivated by that, in this paper, we address the problem of online Bayesian estimation of a categorical distribution ($\theta$) under $\epsilon$-LDP, while at the same time choosing the randomization mechanism adaptively so that the utility is improved continually in time.
@@
Figure 1: AdOBEst-LDP: A framework for Adaptive and Online Bayesian Estimation of categorical distributions with Local Differential Privacy.
This paper presents AdOBEst-LDP: a new methodological framework for Adaptive Online Bayesian Frequency Estimation with LDP. A flowchart diagram of AdOBEst-LDP is given in Figure 1 to expose the reader to the main idea of the framework. The main idea of AdOBEst-LDP is to collect future privatized categorical data with high estimation utility based on the knowledge extracted from the previously collected privatized categorical data. To achieve this goal, AdOBEst-LDP continually adapts its randomized response mechanism to the estimation of $\theta$.
The development of AdOBEst-LDP offers three main contributions to the LDP literature.
@@
  • A new randomized response mechanism: AdOBEst-LDP uses a new adaptive Randomly Restricted Randomized Response mechanism (RRRR) to produce randomized responses under $\epsilon$-LDP. RRRR is a generalization of the standard randomized response mechanism in that it restricts the response to a subset of categories. This subset is selected such that the sensitive information $X$ of the next individual is likely contained in that subset. To ensure this, the subset selection step uses two inputs: (i) a sample for $\theta$ drawn from the posterior distribution of $\theta$ conditional on the past data, (ii) a utility function that scores the informativeness of the randomized response obtained from RRRR when it is run with a given subset. To that end, we propose several utility functions to score the informativeness of the randomized response. The utility functions are based on well-known tools and metrics from probability and statistics, such as Fisher information (Alparslan and Yıldırım, , 2022; Lopuhaä-Zwakenberg et al., , 2022; Steinberger, , 2024; Yıldırım, , 2024), entropy, total variation distance, expected squared error, and probability of honest response, i.e., $Y=X$. We provide some insight into those utility functions both theoretically and numerically. Moreover, we also provide a computational complexity analysis for the proposed utility functions.
@@
  • Posterior sampling: We equip AdOBEst-LDP with a scalable posterior sampling method for parameter estimation. Bayesian estimation is a natural choice for inference when the data is corrupted or censored (Kim et al., , 2011; Liu et al., , 2022; Lone et al., , 2024) and such modification can be statistically modeled. In differential privacy settings, too, Bayesian inference is widely employed (Williams and Mcsherry, , 2010; Karwa et al., , 2014; Foulds et al., , 2016; Alparslan and Yıldırım, , 2022) when the input data is shared with privacy-preserving noise. Standard MCMC methods, such as Gibbs sampling, have a computation complexity quadratic in the number of individuals whose data have been collected. As a remedy to this, similar to Mazumdar et al., (2020), we propose a stochastic gradient Langevin dynamics (SGLD)-based algorithm to obtain approximate posterior samples (Welling and Teh, , 2011). By working on subsets of data, SGLD scales in time.
@@
    The numerical experiments show that AdOBEst-LDP outperforms its non-adaptive counterpart when run with SGLD for posterior sampling. The results also suggest that the utility functions considered in this paper are robust and perform well. The MATLAB code at https://github.com/soneraydin/AdOBEst_LDP can be used to reproduce the results obtained in this paper.
@@
  • Convergence results: Finally, we provide a theoretical analysis of AdOBEst-LDP. We prove two main results:
@@
      If posterior sampling is performed exactly, the expected frequency of the algorithm choosing the best subset (according to the utility function) tends to $1$ as $n$ goes to $\infty$.
    The theoretical results require fairly weak, realistic, and verifiable assumptions.
@@ 2 Related Literature
Frequency estimation under the LDP setting has been an increasingly popular research area in recent years. Along with its basic application (estimation of discrete probabilities from locally privatized data), it is also used for a wide range of other estimation and learning purposes such as estimation of confidence intervals and confidence sets for a population mean (Waudby-Smith et al., , 2023), estimation or identification of heavy hitters (Zhu et al., , 2024) (Wang et al., 2024b, ) (Jia and Gong, , 2019), estimation of quantiles (Cormode and Bharadwaj, , 2022), frequent itemset mining (Zhao et al., , 2023), estimation of degree distribution in social networks (Wang et al., 2024a, ), distributed training of graph neural networks with categorical features and labels (Bhaila et al., , 2024). The methods that are proposed for $\epsilon$-LDP frequency estimation also form the basis of more complex inferential tasks (with some modifications on these methods), such as the release of ‘marginals’ (contingency tables) between multiple categorical features and their correlations, as in the work of Cormode et al., (2018).
@@
AdOBEst-LDP employs RRRR as its randomized mechanism to produce randomized responses. RRRR is a modified version of the Standard Randomized Response mechanism (SRR) (also known as generalized randomized response, $k$-randomized response, and direct encoding in the literature.) Given $X$ as its input, SRR outputs $X$ with probability $\frac{e^{\epsilon}}{e^{\epsilon}+K-1}$, otherwise outputs one of the other categories at random. This is a well-studied mechanism in the DP literature, and the statistical properties of its basic version (such as its estimation variance) can be found in the works by (Wang et al., , 2017) and (Wang et al., , 2020). When $K$ is large, the utility of SRR can be too low. RRRR in AdOBEst-LDP is designed to circumvent this problem by constraining its output to a subset of categories. Unlike SRR, the perturbation probability of responses in our algorithm changes adaptively, depending on the cardinality of the selected subset of categories (which we explain in detail in Section 4) for the privatization of $X$, and the cardinality of its complementary set.
@@
The use of information metrics as utility functions in LDP protocols has been an active line of research in recent years. In the work of Kairouz et al., (2016), information metrics like $f$-divergence and mutual information are used for selecting optimal LDP protocols. In the same vein, Steinberger, (2024) uses Fisher Information as the utility metric for finding a nearly optimal LDP protocol for the frequency estimation problem, and Lopuhaä-Zwakenberg et al., (2022) uses it for comparing the utility of various LDP protocols for frequency estimation and finding the optimal one. In these works, the mentioned information metrics are used statically, i.e., to choose a protocol once and for all, for a given estimation task. The approaches in these works suffer from computational complexity for large values of $K$ because the search space for optimal protocols there grows in the order of $2^{K}$. In some other works, such as Wang et al., (2016), a randomly sampled subset of size $k\leq K$ is used to improve the efficiency of this task, where the optimal $k$ is determined by maximizing the mutual information between real data and the privatized data. However, this approach is also static as the optimal subset size $k$ is selected only once, and the optimization procedure only determines $k$ and not the subset itself. Unlike those static approaches, AdOBEst-LDP dynamically uses the information metric (such as the Fisher Information matrix and the other alternatives in Section 4.3) to select the optimal subset at each time step. In addition, in the subset selection step of AdOBEst-LDP, only $K$ candidate subsets are compared in terms of their utilities at each iteration, enabling computational tractability. This way of tackling the problem requires computing the given information metric for only $K$ times at each iteration. We will provide further details of this approach in Section 4.3 and provide a computational complexity analysis in Section 4.4.
@@
Another use of the Fisher Information in the LDP literature is for bounding the estimation error for a given LDP protocol. For example, Barnes et al., (2020) uses Fisher Information inside van Trees inequality, the Bayesian version of the Cramér-Rao bound (Gill and Levit, , 1995), for bounding the estimation error of various LDP protocols for Gaussian mean estimation and frequency estimation. Again, their work provides rules for choosing optimal protocols for a given $\epsilon$ in a static way. As a similar example, Acharya et al., (2023) derives a general information contraction bound for parameter estimation problems under LDP and shows its relation to van Trees inequality as its special case. To our knowledge, our approach is the first one that adaptively uses a utility metric to dynamically update the inner workings of an LDP protocol for estimating categorical distributions.
@@
The idea of building adaptive mechanisms for improved estimation under the LDP has been studied in the literature, although the focus and methodology of those works differ from ours. For example, Joseph et al., (2019) proposed a two-step adaptive method to estimate the unknown mean parameter of data from Gaussian distribution. In this method, the users are split into two groups, an initial mean estimate is obtained from the perturbed data of the first group, and the data from the second group is transformed adaptively according to that initial estimate. Similarly, Wei et al., (2024) proposed another two-step adaptive method for the mean estimation problem, in which the aggregator first computes a rough distribution estimate from the noisy data of a small sample of users, which is then used for adjusting the amount of perturbation for the data of remaining users. While Joseph et al., (2019); Wei et al., (2024) consider a two-stage method, AdOBEst-LDP seeks to adapt continually by updating its LDP mechanism each time an individual’s information is collected. Similar to our work, Yıldırım, (2024) has recently proposed an adaptive LDP mechanism for online parameter estimation for continuous distributions. The LDP mechanism of Yıldırım, (2024) contains a truncation step with boundaries adapted to the estimate from the past data according to a utility function based on the Fisher information. Unfortunately, the parameter estimation step of Yıldırım, (2024) does not scale in time. Differently from Yıldırım, (2024), AdOBEst-LDP focuses on categorical distributions, considers several other utility functions to update its LDP mechanism, employs a scalable parameter estimation step, and its performance is backed up with theoretical results.